OpenStudy (anonymous):

Solve : 1) $2000(1+.08/4)^4(18)

jimthompson5910 (jim_thompson5910):

2000(1+.08/4)^4(18) 2000(1+0.02)^72 2000(1.02)^72 2000(4.161140375) 8,322.28075 8,322.28 ------------------------------ So \[\Large 2000\left(1+\frac{0.08}{4}\right)^{4(18)} = 8,322.28\]

OpenStudy (anonymous):

Thanks! I went back and did it and that's what I got , miscalculated it . @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok great, glad you figured out where you went wrong

OpenStudy (anonymous):

8000=22000(1-.0125/12)12t A=1500e.075(5)

OpenStudy (anonymous):

@jim_thompson5910 ^^

jimthompson5910 (jim_thompson5910):

ok so let's solve 8000=22000(1-.0125/12)^(12t) for t first

jimthompson5910 (jim_thompson5910):

wait a second, is that really 22,000 or is it supposed to be 2,000?

OpenStudy (anonymous):

It's correct.

jimthompson5910 (jim_thompson5910):

so it's 22,000 ?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

ok thx

jimthompson5910 (jim_thompson5910):

8000=22000(1-.0125/12)^(12t) 8000/22000 = (1-.0125/12)^(12t) 0.363636 = (1-.0125/12)^(12t) ln(0.363636) = ln((1-.0125/12)^(12t)) -1.01160191 = 12t*ln(1-.0125/12) -1.01160191 = 12t*(-0.0010422) -1.01160191/(-0.0010422) = 12t 970.640865 = 12t 12t = 970.640865 t = 970.640865/12 t = 80.88673875

jimthompson5910 (jim_thompson5910):

For the last one A=1500e^(.075(5)) A=1500e^(0.375) A=1500(1.454991) A=2182.4865 A=2182.49

OpenStudy (anonymous):

So the question says: You buy a new car for $22,000. The value decreases by 12.5% each year. When will the car have a value of $8000? What's the answer to that? I'm lost.

jimthompson5910 (jim_thompson5910):

oh it would just be 8000=22000(1-0.125)^t

jimthompson5910 (jim_thompson5910):

solve that for t

OpenStudy (anonymous):

Okay good so that was the answer I got for the third one the second time around, thanks!

OpenStudy (anonymous):

So how would you solve for t? I was stuck on that.And so then the equation isn't setup like I originally had it?

jimthompson5910 (jim_thompson5910):

follow the same idea to solve 8000=22000(1-0.125)^t

jimthompson5910 (jim_thompson5910):

so you would divide both sides by 22000 then take the Ln of both sides that will allow you to pull down the exponent and solve for t

OpenStudy (anonymous):

Can you show me in equation format?

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

8000 = 22000(1-0.125)^t 8000/22000 = (1-0.125)^t 0.363636 = (1-0.125)^t ln(0.363636) = ln((1-0.125)^t) ln(0.363636) = t*ln(0.875) -1.011599 = t*ln(0.875) -1.011599 = t*(-0.133531) -1.011599/(-0.133531) = t 7.575753 = t t = 7.575753

jimthompson5910 (jim_thompson5910):

So it will take roughly 7 and a half years for the car to go from $22,000 in value to $8,000 in value

OpenStudy (anonymous):

Thank you so much for your help. :)

jimthompson5910 (jim_thompson5910):

np